OpenStudy (anonymous):

What are the next three terms in the sequences below? 2, 5, 7, 12, ___, ___, ___ .

OpenStudy (cwtan):

Did u type wrong? i should be 1,2,3,5,7,12 ?

OpenStudy (anonymous):

NO ITS RIGHT

OpenStudy (cwtan):

2,5,7,12,x,y,z 2+5=7 5+7=12 7+12=x 12+x=y x+y=z

OpenStudy (anonymous):

so whats the answer 2, 5, 7, 12, ___, ___, ___ .

OpenStudy (cwtan):

19,31,50

OpenStudy (anonymous):

thx